# Weak Quasistability and Rajchman measures

### v1

@@ 5. Rajchman Measures and Weak Stability
The term comes from the pioneering works of Rajchman on this class of measures (e.g., [18, $\S 5$]). Since Borel measures over a compact metric space are regular (see, e.g., [4, Corollary 10.6]), Rajchman measures are regular because they are finite (and so Borel) acting over the compact set $\mathbb{T}\kern 0.5pt$}. Take the measure $\eta\!:\Re_{[\kern 1.0pt0,1)}\!\to{\mathbb{R}\kern 0.5pt}$} induced by the measure $\mu\!:\Re_{\mathbb{T}\kern 0.5pt}\!\to{\mathbb{R}\kern 0.5pt}$} via a function $\gamma\!:{[\kern 1.0pt0,1)}\kern-1.0pt\to\!{\mathbb{T}\kern 0.5pt}$}, defined by $\eta(A)=\mu(\gamma(A)\kern 0.5pt)$} for every $A\in\Re_{[\kern 1.0pt0,1)}$}. Here the function $\gamma$ is given by $\gamma(\alpha)={e^{2\pi i\alpha}}$ for every $\alpha\in[\kern 1.0pt0,1)$}, which is measurable, invertible, with a measurable inverse. The above integral sets the Fourier transform $\widetilde{\eta}$ of $\eta$, namely,
@@
For a survey on Rajchman measures, see [14] (also [10, Chapter IX]). The basic properties required here are listed below (see, e.g., [14, p.364 and Theorem 3.4]).
Every absolutely continuous finite measure is Rajchman. Every Rajchman measure is continuous. There exist singular Rajchman measures.
So every singular Rajchman measure is singular-continuous. That every absolutely continuous finite measure is Rajchman is readily verified by using the Riemann–Lebesgue Lemma and the Radon–Nikodym Theorem. That every Rajchman measure is continuous is a direct consequence of Neder’s answer [17, Section 1] to a question posed by Riesz [19, p.315].
The Cantor–Lebesgue measure $\eta_{\kern 1.0pt[\kern 1.0pt0,1)}$ on $\Re_{[\kern 1.0pt0,1)}$ is a classical example of a singular-continuous measure on $\Re_{[\kern 1.0pt0,1)}.$ This is the Borel-Stieltjes measure on $\Re_{[\kern 1.0pt0,1)}$ generated by the Cantor function associated with the Cantor set in $0,1\kern-1.0pt]$}. The Cantor–Lebesgue measure $\mu_{{}_{{\scriptstyle{\mathbb{T}\kern 0.5pt}}}\!}\kern-1.0pt$ on $\Re_{\mathbb{T}\kern 0.5pt}\kern-1.0pt$ is obtained from $\eta_{\kern 1.0pt[\kern 1.0pt0,1)}$ as $\mu_{{}_{{\scriptstyle{\mathbb{T}\kern 0.5pt}}}\!}(E)\kern-1.0pt=\kern-1.0pt\mu_{[\kern 1.0pt0,1)}\kern-0.5pt\big{(}\gamma^{-1}(E)\big{)}$} for $E\in\Re_{\mathbb{T}\kern 0.5pt}$}, where the transformation $\gamma\!:[\kern 1.0pt0,1)\to\kern-1.0pt{\mathbb{T}\kern 0.5pt}$} was defined above. Thus $\mu_{{}_{{\scriptstyle{\mathbb{T}\kern 0.5pt}}}\!}$ is a singular-continuous measure on $\Re_{\mathbb{T}\kern 0.5pt}.\kern-1.0pt$ This, however, is not a Rajchman measure. Indeed, it can be verified that $\int_{\mathbb{T}\kern 0.5pt}\kern-1.0ptz^{k}d\mu_{{}_{{\scriptstyle{\mathbb{T}\kern 0.5pt}}}\!}=\!\int_{\mathbb{T}\kern 0.5pt}\kern-1.0ptz^{3k}d\mu_{{}_{{\scriptstyle{\mathbb{T}\kern 0.5pt}}}}\!$} (i.e., $\tilde{\eta}_{{}_{\,{\scriptstyle{\mathbb{T}\kern 0.5pt}}}\!}(k)=\tilde{\eta}_{{}_{\,{\scriptstyle{\mathbb{T}\kern 0.5pt}}}\!}(3k)\kern 0.5pt$}) for all $k\in{\mathbb{Z}\kern 0.5pt}$}. The first example of a singular Rajchman measure is due to Menchoff [16, Lemma p.433 and Theorem p.435], which is a modification of the Cantor–Lebesgue measure on $\Re_{\mathbb{T}\kern 0.5pt}.$
@@
$\langle U_{\varphi,\mu}^{k}\psi\,;\phi\rangle=\!\int_{\mathbb{T}\kern 0.5pt}z^{k}\psi\kern 1.0pt\overline{\phi}\,d\mu\quad\hbox{so that}\quad\langle U_{\varphi,\mu}^{k}1\,;1\rangle=\!\int_{\mathbb{T}\kern 0.5pt}z^{k}d\mu,$
@@
Remark 5.1. A standard application of the Stone–Weierstrass Theorem. Let $P({\mathbb{T}\kern 0.5pt})$ be the set of all polynomials $q(\;\;,\;\;)\!:\!{\mathbb{T}\kern 0.5pt}\!\times\!{\mathbb{T}\kern 0.5pt}\kern-1.0pt\to{\mathbb{C}\kern 0.5pt}$} in $z$ and $\overline{z}$ (of the form $\sum_{k,\ell=0}^{N}\alpha_{k,\ell}\,z^{k}z^{-\ell}$} with $z^{k}\kern-1.0pt,{\overline{z}}^{\,\ell}\kern-1.0pt\in\kern-1.0pt{\mathbb{T}\kern 0.5pt}$} and ${\alpha_{k,\ell}\kern-1.0pt\in\kern-1.0pt{\mathbb{C}\kern 0.5pt}}.)$ This is a classical density result:
@@
for every $E$ in $\Re_{T}$ and every $k$ in $\mathbb{Z}\kern 0.5pt$}. There are two ways to conclude the proof. Conclusion 1. As $\eta_{g}\kern-0.5pt$ is Rajchman and $g^{-1}\kern-1.0pt\in L^{1}({\mathbb{T}\kern 0.5pt},\mu)$}, $\int_{\mathbb{T}\kern 0.5pt}z^{k}g^{-1}d\kern 1.0pt\eta_{g}\kern-1.0pt\to 0$} as $|k|\to\infty$} because $\rm(a)\Longrightarrow(c)$}. So $\mu$ is Rajchman by $(\star)$ with $E=\kern-1.0pt{\mathbb{T}\kern 0.5pt}$}. Thus (a) holds true. Conclusion 2. For k=0{k=0}, $(\star)$ says that $\mu(E)=\int_{E}d\mu=\!\int_{E}g^{-1}d\kern 1.0pt\eta_{g}$} for every $E\in\Re_{\mathbb{T}\kern 0.5pt}$}, and so $\mu$ is absolutely continuous with respect to $\eta_{g}$ (i.e., ${\mu\kern-1.0pt\ll\kern-1.0pt\eta_{g}}).$ Since $\eta_{g}\kern-1.0pt$ is Rajchman, it follows that $\mu$ must be Rajchman. (This is a result due to Milicer-Grużewska (cf. [14, p.365 (2.1)]). Thus, again, (a) holds true.
@@ 6. Rajchman Measures and Trigonometric Basis
(a1) With ${\bf e}_{k}(z)=z^{k}\in{\mathbb{T}\kern 0.5pt}$} for every $z\in{\mathbb{T}\kern 0.5pt}$} and $\{{\bf e}_{k}\}$ being an orthonormal set, we get $1=\|{\bf e}_{k}\|^{2}\!=\!\int_{\mathbb{T}\kern 0.5pt}|{\bf e}_{k}(z)|^{2}d\mu(z)=\!\int_{\mathbb{T}\kern 0.5pt}|z^{k}|^{2}d\mu(z)=\!\int_{\mathbb{T}\kern 0.5pt}d\mu=\mu({\mathbb{T}\kern 0.5pt})$ for every $k$ in $\mathbb{Z}\kern 0.5pt$}, so that $\mu$ is a probability measure. (a2) Using the preceding argument, we can verify that $\int_{\mathbb{T}\kern 0.5pt}z^{n}d\mu\to 0$} as $n\to\infty$}. In fact, let $U_{\varphi,\mu}\kern-1.0pt$ be the position operator on $L^{2}({{\mathbb{T}\kern 0.5pt},\mu})$. If $\{z^{k}\}$ is an orthonormal basis for $L^{2}({\mathbb{T}\kern 0.5pt},\mu)$, then $U_{\varphi,\mu}$ is a bilateral shift (because it shifts this orthonormal basis; i.e., $U_{\varphi,\mu}{\bf e}_{k}={\bf e}_{k+1}$}), and therefore, like every bilateral shift, it is weakly stable, so that ${\langle U_{\varphi,\mu}^{n}1\,;1\rangle}={\int_{\mathbb{T}\kern 0.5pt}z^{n}d\mu\to 0}$ as $n\to\infty$}. This implies $\int_{\mathbb{T}\kern 0.5pt}z^{k}d\mu\to 0$} as $|k|\to\infty$}. Since $\mu$ is finite by (a), it is a Rajchman measure. (b) The set $\{z^{k}\}$ is made up of unit vectors because $\mu$ is a probability measure. Let $\mu$ be singular with respect to the normalised Lebesgue measure $\lambda$ on $\Re_{\mathbb{T}\kern 0.5pt}$, with support $\mu]=C\subseteq\{z\in{\mathbb{T}\kern 0.5pt}\!:z=e^{2\pi i\alpha}\;\hbox{for every}\;\alpha\in[\kern 1.0pt0,\frac{1}{4}]\}$}. Thus there exists an $\Re_{\mathbb{T}\kern 0.5pt}$-measurable partition $\{B,C\}$ of $\mathbb{T}\kern 0.5pt$} such that $\mu(B)=\lambda(C)=0$} and ${\lambda({\mathbb{T}\kern 0.5pt})=\lambda(B)}={\mu(C)=\mu({\mathbb{T}\kern 0.5pt})=1}.$ Therefore,
@@
with $\eta\!:\Re_{[\kern 1.0pt0,1)}\to[\kern 1.0pt0,1]$} given by $\eta(A)=\mu(\gamma(A)\kern 0.5pt)$} for every set $A\in\Re_{[\kern 1.0pt0,1)}$}, where $\gamma\!:[\kern 1.0pt0,1)\to{\mathbb{T}\kern 0.5pt}$} is the homeomorphism (thus a measurable transformation) of $\kern 1.0pt0,1)$} (equipped with the usual metric in $\mathbb{R}\kern 0.5pt$}) onto $\mathbb{T}\kern 0.5pt$} (equipped with the arc-length metric in $\mathbb{T}\kern 0.5pt$}) defined by $\gamma(\alpha)\kern-1.0pt=\kern-1.0pte^{2\pi i\alpha}$} for every $\alpha\kern-1.0pt\in\kern-1.0pt[\kern 1.0pt0,1)$}. The support of $\eta$ is $\eta]=\gamma^{-1}(\kern 0.5pt[\mu]\kern 0.5pt)$} $\gamma^{-1}(C)\subseteq[\kern 1.0pt0,\frac{1}{4}]$}. If $\mu$ is continuous with respect to $\lambda$, then $\eta$ is continuous with respect to the restriction to $\kern 1.0pt0,1)$} of the Lebesgue measure on $\Re$, and therefore singletons in $\kern 1.0pt0,\frac{1}{4}]$} have $\eta$-measure zero. Hence, since $\eta$ is concentrated on $\kern 1.0pt0,\frac{1}{4}]$},
@@
Remark 6.2. Just orthogonality is lost. However, the standard application of the Stone–Weierstrass Theorem summarised in Remark 5.1 says that, for any finite measure $\mu$ on $\Re_{\mathbb{T}\kern 0.5pt}$, functions on $L^{p}({\mathbb{T}\kern 0.5pt},\mu)$ are approximated by the trigonometric polynomials in $P({\mathbb{T}\kern 0.5pt}).$ In other words, using the argument in the proof of Proposition 5.3 (for $\rm(a)\Longrightarrow(b)$}, and so for ${\rm(a)\Longrightarrow(e)}\kern 0.5pt)$, for every $p\kern-1.0pt\geq\kern-1.0pt1$}, $\bigvee_{k\in{\mathbb{Z}\kern 0.5pt}}\!\{z^{k}\}\!=L^{p}({\mathbb{T}\kern 0.5pt},\mu)$}, where $\bigvee_{k\in{\mathbb{Z}\kern 0.5pt}}\!\{z^{k}\}$ is the closure of the linear span of $\{z^{k}\}$ (see, e.g., [5, Theorem 7.2] for the case of Lebesgue measure on $\Re_{\mathbb{T}\kern 0.5pt}$).

### v2

@@ 5. Rajchman Measures and Weak Stability
The term comes from the pioneering works of Rajchman on this class of measures (e.g., [19, $\S 5$]). Since Borel measures over a compact metric space are regular (see, e.g., [4, Corollary 10.6]), Rajchman measures are regular because they are finite (and so Borel) acting over the compact set $\mathbb{T}\kern 0.5pt$}. Take the measure $\eta\!:\Re_{[\kern 1.0pt0,1)}\!\to{\mathbb{R}\kern 0.5pt}$} induced by the measure $\mu\!:\Re_{\mathbb{T}\kern 0.5pt}\!\to{\mathbb{R}\kern 0.5pt}$} via a function $\gamma\!:{[\kern 1.0pt0,1)}\kern-1.0pt\to\!{\mathbb{T}\kern 0.5pt}$}, defined by $\eta(A)=\mu(\gamma(A)\kern 0.5pt)$} for every $A\in\Re_{[\kern 1.0pt0,1)}$}. Here the function $\gamma$ is given by $\gamma(\alpha)={e^{2\pi i\alpha}}$ for every $\alpha\in[\kern 1.0pt0,1)$}, which is measurable, invertible, with a measurable inverse. The above integral sets the Fourier transform $\widetilde{\eta}$ of $\eta$, namely,
@@
For a survey on Rajchman measures, see [15] (also [10, Chapter IX]). The basic properties required here are listed below (see, e.g., [15, p.364 and Theorem 3.4]).
Every absolutely continuous finite measure is Rajchman (in particular, the normalised Lebesgue measure on $\Re_{\mathbb{T}\kern 0.5pt}$ is Rajchman). Every Rajchman measure is continuous. There exist singular Rajchman measures (and so every singular Rajchman measure is singular-continuous).
That every absolutely continuous finite measure is Rajchman is readily verified by the Riemann–Lebesgue Lemma and the Radon–Nikodym Theorem. That every Rajchman measure is continuous is a direct consequence of Neder’s answer [18, Section 1] to a question posed by Riesz [20, p.315].
The Cantor–Lebesgue measure $\eta_{\kern 1.0pt[\kern 1.0pt0,1)}$ on $\Re_{[\kern 1.0pt0,1)}$ is a classical example of a singular-continuous measure on $\Re_{[\kern 1.0pt0,1)}.$ This is the Borel-Stieltjes measure on $\Re_{[\kern 1.0pt0,1)}$ generated by the Cantor function associated with the Cantor set in $0,1\kern-1.0pt]$}. The Cantor–Lebesgue measure $\mu_{{}_{{\scriptstyle{\mathbb{T}\kern 0.5pt}}}\!}\kern-1.0pt$ on $\Re_{\mathbb{T}\kern 0.5pt}\kern-1.0pt$ is obtained from $\eta_{\kern 1.0pt[\kern 1.0pt0,1)}$ as $\mu_{{}_{{\scriptstyle{\mathbb{T}\kern 0.5pt}}}\!}(E)\kern-1.0pt=\kern-1.0pt\mu_{[\kern 1.0pt0,1)}\kern-0.5pt\big(\gamma^{-1}(E)\big)$} for $E\in\Re_{\mathbb{T}\kern 0.5pt}$}, where the transformation $\gamma\!:[\kern 1.0pt0,1)\to\kern-1.0pt{\mathbb{T}\kern 0.5pt}$} was defined above. Thus $\mu_{{}_{{\scriptstyle{\mathbb{T}\kern 0.5pt}}}\!}$ is a singular-continuous measure on $\Re_{\mathbb{T}\kern 0.5pt}.\kern-1.0pt$ This, however, is not a Rajchman measure. Indeed, it can be verified that $\int_{\mathbb{T}\kern 0.5pt}\kern-1.0ptz^{k}d\mu_{{}_{{\scriptstyle{\mathbb{T}\kern 0.5pt}}}\!}=\!\int_{\mathbb{T}\kern 0.5pt}\kern-1.0ptz^{3k}d\mu_{{}_{{\scriptstyle{\mathbb{T}\kern 0.5pt}}}}\!$} (i.e., $\tilde{\eta}_{{}_{\,{\scriptstyle{\mathbb{T}\kern 0.5pt}}}\!}(k)=\tilde{\eta}_{{}_{\,{\scriptstyle{\mathbb{T}\kern 0.5pt}}}\!}(3k)\kern 0.5pt$}) for all $k\in{\mathbb{Z}\kern 0.5pt}$}. The first example of a singular Rajchman measure is due to Menchoff [17, Lemma p.433 and Theorem p.435], which is a modification of the Cantor–Lebesgue measure on $\Re_{\mathbb{T}\kern 0.5pt}.$
@@
$\langle U_{\varphi,\mu}^{k}\psi\,;\phi\rangle=\!\int_{\mathbb{T}\kern 0.5pt}z^{k}\psi\kern 1.0pt\overline{\phi}\,d\mu\qquad\hbox{so that}\qquad\langle U_{\varphi,\mu}^{k}1\,;1\rangle=\!\int_{\mathbb{T}\kern 0.5pt}z^{k}d\mu,$
@@
Remark 5.1. A standard application of the Stone–Weierstrass Theorem.
Let $P({\mathbb{T}\kern 0.5pt})$ denote the set of all polynomials $q(\;\;,\;\;)\!:\!{\mathbb{T}\kern 0.5pt}\!\times\!{\mathbb{T}\kern 0.5pt}\kern-2.0pt\to{\mathbb{C}\kern 0.5pt}$} in $z$ and $\overline{z}$ (of the form $\sum_{k,\ell=0}^{N}\alpha_{k,\ell}\,z^{k}z^{-\ell}$} with $z^{k}\kern-1.0pt,{\overline{z}}^{\,\ell}\kern-1.0pt\in\kern-1.0pt{\mathbb{T}\kern 0.5pt}$} and ${\alpha_{k,\ell}\kern-1.0pt\in\kern-1.0pt{\mathbb{C}\kern 0.5pt}}.)$ This is a classical density result:
@@
Remark 5.2. A basic change of variables for complex measurable functions.
@@
for every $E$ in $\Re_{T}$ and every $k$ in $\mathbb{Z}\kern 0.5pt$}. There are two ways to conclude the proof.
Conclusion 1. As $\eta_{g}\kern-0.5pt$ is Rajchman and $g^{-1}\kern-1.0pt\in L^{1}({\mathbb{T}\kern 0.5pt},\mu)$}, $\int_{\mathbb{T}\kern 0.5pt}z^{k}g^{-1}d\kern 1.0pt\eta_{g}\kern-1.0pt\to 0$} as $|k|\to\infty$} because $\rm(a)\Longrightarrow(c)$}. So $\mu$ is Rajchman by $(\star)$ with $E=\kern-1.0pt{\mathbb{T}\kern 0.5pt}$}. Thus (a) holds true.
Conclusion 2. For k=0{k=0}, $(\star)$ says that $\mu(E)=\int_{E}d\mu=\!\int_{E}g^{-1}d\kern 1.0pt\eta_{g}$} for every $E\in\Re_{\mathbb{T}\kern 0.5pt}$}, and so $\mu$ is absolutely continuous with respect to $\eta_{g}$ (i.e., ${\mu\kern-1.0pt\ll\kern-1.0pt\eta_{g}}).$ Since $\eta_{g}\kern-1.0pt$ is Rajchman, it follows that $\mu$ must be Rajchman. (This is a result due to Milicer-Grużewska (cf. [15, p.365 (2.1)]). Thus, again, (a) holds true.
@@ 6. Rajchman Measures and Trigonometric Basis
(a1) With ${\bf e}_{k}(z)=z^{k}\in{\mathbb{T}\kern 0.5pt}$} for every $z\in{\mathbb{T}\kern 0.5pt}$} and $\{{\bf e}_{k}\}$ being an orthonormal set, we get $1=\|{\bf e}_{k}\|^{2}\!=\!\int_{\mathbb{T}\kern 0.5pt}|{\bf e}_{k}(z)|^{2}d\mu(z)=\!\int_{\mathbb{T}\kern 0.5pt}|z^{k}|^{2}d\mu(z)=\!\int_{\mathbb{T}\kern 0.5pt}d\mu=\mu({\mathbb{T}\kern 0.5pt})$ for every $k$ in $\mathbb{Z}\kern 0.5pt$}, so that $\mu$ is a probability measure.
(a2) Using the preceding argument, we can verify that $\int_{\mathbb{T}\kern 0.5pt}z^{n}d\mu\to 0$} as $n\to\infty$}. In fact, let $U_{\varphi,\mu}\kern-1.0pt$ be the position operator on $L^{2}({{\mathbb{T}\kern 0.5pt},\mu})$. If $\{z^{k}\}$ is an orthonormal basis for $L^{2}({\mathbb{T}\kern 0.5pt},\mu)$, then $U_{\varphi,\mu}$ is a bilateral shift (because it shifts this orthonormal basis; i.e., $U_{\varphi,\mu}{\bf e}_{k}={\bf e}_{k+1}$}). Thus, like every bilateral shift, it is weakly stable, so that ${\langle U_{\varphi,\mu}^{n}1\,;1\rangle}={\int_{\mathbb{T}\kern 0.5pt}z^{n}d\mu\to 0}$ as $n\to\infty$}. This implies $\int_{\mathbb{T}\kern 0.5pt}z^{k}d\mu\to 0$} as $|k|\to\infty$}. Since $\mu$ is finite by (a), it is a Rajchman measure.
(b) The set $\{z^{k}\}$ is made up of unit vectors because $\mu$ is a probability measure. Let $\mu$ be singular with respect to the normalised Lebesgue measure $\lambda$ on $\Re_{\mathbb{T}\kern 0.5pt}$, with support $\mu]=C\subseteq\{z\in{\mathbb{T}\kern 0.5pt}\!:z=e^{2\pi i\alpha}\;\hbox{for every}\;\alpha\in[\kern 1.0pt0,\frac{1}{4}]\}$}. So there exists an $\Re_{\mathbb{T}\kern 0.5pt}$-measurable partition $\{B,C\}$ of $\mathbb{T}\kern 0.5pt$} such that $\mu(B)=\lambda(C)=0$} and ${\lambda({\mathbb{T}\kern 0.5pt})=\lambda(B)}={\mu(C)=\mu({\mathbb{T}\kern 0.5pt})=1}.$ Hence,
@@
with $\eta\!:\Re_{[\kern 1.0pt0,1)}\to[\kern 1.0pt0,1]$} given by $\eta(A)=\mu(\gamma(A)\kern 0.5pt)$} for every set $A\in\Re_{[\kern 1.0pt0,1)}$}, where $\gamma\!:[\kern 1.0pt0,1)\to{\mathbb{T}\kern 0.5pt}$} is the homeomorphism (thus a measurable transformation) of $\kern 1.0pt0,1)$} (equipped with the usual metric in $\mathbb{R}\kern 0.5pt$}) onto $\mathbb{T}\kern 0.5pt$} (equipped with the arc-length metric in $\mathbb{T}\kern 0.5pt$}) defined by $\gamma(\alpha)\kern-1.0pt=\kern-1.0pte^{2\pi i\alpha}$} for every $\alpha\kern-1.0pt\in\kern-1.0pt[\kern 1.0pt0,1)$}. The support of $\eta$ is $\eta]=\gamma^{-1}(\kern 0.5pt[\mu]\kern 0.5pt)$} $\gamma^{-1}(C)\subseteq[\kern 1.0pt0,\frac{1}{4}]$}. If $\mu$ is continuous with respect to $\lambda$, then $\eta$ is continuous with respect to the restriction to $\kern 1.0pt0,1)$} of the Lebesgue measure on $\Re$, and therefore singletons in $\kern 1.0pt0,\frac{1}{4}]$} have $\eta$-measure zero. Therefore, since $\eta$ is concentrated on $\kern 1.0pt0,\frac{1}{4}]$},
@@
Remark 6.2. Just orthogonality is lost.
However, the standard application of the Stone–Weierstrass Theorem summarised in Remark 5.1 says that, for any finite measure $\mu$ on $\Re_{\mathbb{T}\kern 0.5pt}$, functions on $L^{p}({\mathbb{T}\kern 0.5pt},\mu)$ are approximated by the trigonometric polynomials in $P({\mathbb{T}\kern 0.5pt}).$ In other words, using the argument in the proof of Proposition 5.3 (for $\rm(a)\Longrightarrow(b)$}, and so for ${\rm(a)\Longrightarrow(e)}\kern 0.5pt)$, for every $p\kern-1.0pt\geq\kern-1.0pt1$}, $\bigvee_{k\in{\mathbb{Z}\kern 0.5pt}}\!\{z^{k}\}\!=L^{p}({\mathbb{T}\kern 0.5pt},\mu)$}, where $\bigvee_{k\in{\mathbb{Z}\kern 0.5pt}}\!\{z^{k}\}$ is the closure of the linear span of $\{z^{k}\}$ (see, e.g., [5, Theorem 7.2] for the case of Lebesgue measure on $\Re_{\mathbb{T}\kern 0.5pt}$).